Circular No. 6658

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 B. Schaefer and M. Schaefer, Yale University; P. Smith and C. Mackey,

Kitt Peak National Observatory; E. Wilcots, D. J. Pisano, and C. Howk, University of Wisconsin; E. Costa, M. Feroci, and L. Piro, Istituto di Astrofisica Spaziale, Frascati; M. Tavani, Columbia University and Istituto di Fisica Cosmica e Tecnologie Relative, Milan; F. Frontera, Università di Ferrara; and E. Palazzi and L. Nicastro, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna, report on observations with the 3.5-m WIYN telescope on May 9.203, 10.142, and 12.135 UT: “There are seven sources inside the 50′′ BeppoSAX/NFI error box (IAUC 6656). The R magnitudes of these sources were obtained with respect to a star at α = 6^{h}53^{m}48^{s}.5, δ = +79^{o}16′32′′.7 (equinox 2000.0), which has R = 19.70 (this magnitude is based on calibration of the filter and detector on previous nights). Following are given the position end figures (accuracy ∼ 0′′.5) for the other six sources, the respective R magnitudes (accurate to typically 0.02 mag for the brighter sources and 0.15 mag for the galaxy), and apparent object types: 38^{s}.9, 15′22′′.0, 17.25, stellar; 36^{s}.1, 15′31′′.1, 18.91, stellar; 49^{s}.4, 16′20′′.0, 21.40–20.30, stellar variable; 54^{s}.8, 16′03′′.7, 21.34, stellar; 41^{s}.3, 15′43′′.5, 20.55, unusual; 45^{s}.2, 15′57′′.0, ∼ 22.7, galaxy. Bond’s variable (see IAUC 6654) is a point source that has brightened from R = 21.40 ± 0.05 on May 9 to 20.30 ± 0.02 on May 10, fading to 20.47 ± 0.03 on May 11. The existence of brightening and fading creates a coincidence in time that can causally connect the variable with the burst, and matches various model predictions. The ‘unusual’ object is an extended source shaped like a ‘Z’, extending 7′′ from tip-to-tip; the morphology of this source is highly unusual, although it could be a superposition of three or more sources.”

S. G. Djorgovski, S. C. Odewahn, R. R. Gal, M. R. Metzger, S. R. Kulkarni, and M. A. Pahre, California Institute of Technology; D. A. Frail, National Radio Astronomy Observatory; and E. Costa, M. Feroci, and the rest of the BeppoSAX team write: “Observations at Palomar indicate that the optical counterpart of GRB 970508 (IAUC 6654, 6655) has began to decline in brightness, as shown by the following Gunn r-band magnitudes: May 9.195 UT, 21.2 ± 0.15; 10.178, 20.18 ± 0.07; 11.198, 20.16 ± 0.06; 12.195, 20.53 ± 0.06. It had apparently gone through a peak brightness some time on May 10, and is now fading. Such lightcurve behavior is expected in some classes of theoretical models.”

1997 May 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green